OpenStudy (anonymous):

56 of what number is 75.6? help please

OpenStudy (jamesj):

56% you mean?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

thts what i ment lol

OpenStudy (jamesj):

56% of a number, x, is \[ \frac{56}{100}x \] Hence you want to find the number x such that \[ \frac{56}{100}x = 75.6 \]

OpenStudy (anonymous):

56% of x is 0.56x. Set this equal to 75.6 and we have: 0.56x = 75.6. Next solve for x: x = 75.6/0.56 = 7560/56 = 135